A rope holds a 10-kg rock at rest on a frictionless inclined plane of 30 degrees. Determine the magnitude of the acceleration of the rock as it slides down theinclined plane if the rope breaks?
Answer #1

accleration = gsin 30 = 9.8*1/2 = 4.9m/s2
Answer #3
G=m*g=10kg * 10m/s^2=100N (force of gravity directed downwards)

R is Reaction force perpendicular to the inclined plane.

The sum of G and R create a force that acts along the inclines plane and causes the tension in the rope T.

(Make a sketch to see it better)

T=G*sin30=100N * 0.5=50N